Questions Quiz 17-6.02m d Question 6 of 15 D Check My Work (1 remaining) Assume that 180-day U.S. securities have a 5.5% annual interest rate, whereas 180-day British securities have an 8.0% annual interest rate. In the spot market, one British pound 3 can be exchanged for $1.7471. If interest rate parity holds, what is the 180-day forward o exchange rate between U.S. dollars and British pounds quoted in American terms? 7. a.E0.579/$ b. $1.333/E Oc. $1.726/E O d. $2.000/E e. E0.750/$ 10 12. 13 15
Answer #1

180 forward rate in terms of US dollars = Spot Rate *(1+US rate*180/365)/(1+British Rate*180/365) =
1.7471*(1+0.055*180/365)/(1+.08*180/365) = 1.726
Option c is correct

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well